OpenStudy (anonymous):

differentiate x^3(1-3x)^(1/4)

OpenStudy (anonymous):

Product rule?

OpenStudy (anonymous):

chain rule is not needed?

OpenStudy (astrophysics):

\[x^3(1-3x)^{1/4}\] yup product rule should do!

OpenStudy (astrophysics):

\[f'g+g'f\]

OpenStudy (astrophysics):

You will have to use chain rule as well inside the product rule

OpenStudy (anonymous):

f=x^3 and g=(1-x)^1/4

OpenStudy (welshfella):

yep

OpenStudy (nincompoop):

you can also use distributive property first and then use sum

OpenStudy (nincompoop):

\(\huge x^n (1 - ax^\frac{p}{q}) = x^n - ax^\frac{pn}{q} \) then differentiate using power rule

OpenStudy (nincompoop):

\(\huge nx^{n-1} - (pn)ax^{\frac{pn}{q}-1} \)